# ESPFormer: Doubly-Stochastic Attention with Expected Sliced Transport Plans

### v1

@@ 4.4 Vision Transformers
To evaluate the generalizability of the models under limited data scenarios, we conducted experiments on the Cats and Dogs dataset [16] using varying fractions of the training data: 1%, 10%, 25%, and 100%. We train a ViT and modify the attention mechanism accordingly. The models compared include ESPFormer, Sinkformer, DiffTransformer, and the standard Transformer. The goal was to assess how well each model performs when faced with progressively larger amounts of data, particularly in resource-constrained settings. Table 4 summarizes the classification accuracy for each model under different data fractions. Our results demonstrate that ESPFormer consistently outperforms the other models across all data fractions, achieving the highest classification accuracy at each level of data availability. With only 1% of the data, ESPFormer achieves 55.66% accuracy, a significant improvement over the Transformer at 49.71%. As the data availability increases, ESPFormer continues to show superior performance. This suggests that ESPFormer is particularly robust in data-scarce environments, showcasing its ability to generalize well even under limited training data.
Table 4: Average and standard deviation (over 3 runs) of ESPFormer’s classification accuracy (%) vs. baselines on the Cats and Dogs dataset under varying data availability.
      Data Fraction       ESPFormer       Sinkformer       DiffTransformer       Transformer
      1%       55.66 $\pm$ 3.95       55.07 $\pm$ 3.34       53.78 $\pm$ 0.28       49.71 $\pm$ 0.31
      10%       71.49 $\pm$ 0.43       69.56 $\pm$ 0.32       67.34 $\pm$ 0.11       57.25 $\pm$ 0.22
      25%       75.40 $\pm$ 0.38       74.56 $\pm$ 0.58       74.86 $\pm$ 0.17       72.25 $\pm$ 0.16
      100%       79.47 $\pm$ 0.12       79.12 $\pm$ 0.17       78.85 $\pm$ 0.11       78.49 $\pm$ 0.09
Impact of Patch Size
Similar to Sander et al. [44] and to analyze the effect of patch size on model performance, we trained Transformer, DiffTransformer, Sinkformer, and ESPFormer on the MNIST dataset [23]. To isolate the impact of the attention mechanism, we used a single-layer self-attention module without feed-forward layers. Figure 4 illustrates the test accuracy as a function of patch size for each model. ESPFormer consistently outperforms competing models, particularly for smaller patch sizes, highlighting its superior ability to capture fine-grained details. As the patch size increases, the performance gap narrows, with all models converging to similar accuracies — likely due to the reduced information content within each individual patch.
@@ 5 Conclusion
We introduced ESPFormer, a novel Transformer architecture that integrates a fast, doubly-stochastic attention mechanism based on Expected Sliced Transport Plans (ESP). By leveraging differentiable soft sorting and a fully parallelizable approach, ESPFormer provides a computationally efficient alternative to iterative Sinkhorn normalization while ensuring a balanced distribution of attention across tokens. Our experiments across diverse domains, including image classification, point cloud processing, sentiment analysis, and neural machine translation, demonstrate that ESPFormer consistently outperforms both classical self-attention and Sinkhorn-based alternatives in terms of accuracy and efficiency. Furthermore, we showed that our proposed attention can be seamlessly integrated into pre-trained Transformers, improving performance even with minimal fine-tuning. The flexibility of ESPFormer also makes it compatible with emerging differential attention mechanisms, e.g., [55], expanding its applicability to a broad range of architectures. These findings highlight the potential of transport-based attention mechanisms as a principled alternative to existing methods, paving the way for future research in efficient and structured attention models.
@@ Acknowledgment
This work was partially supported by NSF CAREER Award #2339898, and the Wellcome Leap ‘Surgery: Access Validate and Expand (SAVE)’ program.
Appendix A Full runtime wall-clock Analysis
@@ C.1 Table of Notations
$\tau$ The "inverse temperature" hyperparameter in ESP
@@ C.2 ModelNet 40
The model architecture consists of two Induced Set Attention Blocks (ISAB) in the encoder, followed by a decoder incorporating a Set Attention Block (SAB) and a Pooling by Multihead Attention (PMA) module. The training procedure employs a batch size of 64 and utilizes the Adam optimizer [19]. The network is trained for 300 epochs, with an initial learning rate of $10^{-3}$, which is reduced by a factor of 10 after 200 epochs.
@@ C.5.1 Cats and Dogs Classification
For this experiment, we use the ViT model [7] with different attention mechanisms. The images are resized to $224\times 224$. We use a ViT architecture with an embedding and MLP dimension of 128, 6 layers, 8 attention heads, and a patch size of 16. For all methods and percentages of data, we train the model for 300 epochs. We use an initial learning rate of $3\times 10^{-5}$. After 250 epochs, the learning rate is reduced by a factor of 10. Training is done using the Adam optimizer [19] and a batch size of 64. Our experimental setup, including the normalizations and data augmentations, are consistent with the Cats and Dogs experiment in Sinkformer [44] . For each percentage of the data, three random seeds are used to initialize and generate new subsets of data. The subsets are consistent across all methods to ensure the same training set is used.

### v2

@@ 4.4 Vision Transformers
To evaluate the generalizability of the models under limited data scenarios, we conducted experiments on the Cats and Dogs dataset (Kaggle, 2013) using varying fractions of the training data: 1%, 10%, 25%, and 100%. We train a ViT and modify the attention mechanism accordingly. The models compared include ESPFormer, Sinkformer, DiffTransformer, and the standard Transformer. The goal was to assess how well each model performs when faced with progressively larger amounts of data, particularly in resource-constrained settings. Table 1 summarizes the classification accuracy for each model under different data fractions. Our results demonstrate that ESPFormer consistently outperforms the other models across all data fractions, achieving the highest classification accuracy at each level of data availability. With only 1% of the data, ESPFormer achieves 55.66% accuracy, a significant improvement over the Transformer at 49.71%. As the data availability increases, ESPFormer continues to show superior performance. This suggests that ESPFormer is particularly robust in data-scarce environments, showcasing its ability to generalize well even under limited training data.
Transition to Hard Sorting
During training, the use of soft sorting—which depends on a temperature parameter—results in approximate doubly stochastic matrices (DSMs). However, this approximation can be addressed by gradually annealing the temperature during training, effectively transitioning toward hard sorting. Consequently, at inference time, soft sorting can be replaced with hard sorting to yield exact DSMs. This not only ensures exact computations but also reduces the inference time from $O(N^{2})$ to $O(LN\log N)$, where $L$ is the number of slices.
To further clarify this point, we conducted additional experiments the Cats vs. Dogs dataset using the same fractions of the training data to evaluate the effect of hard sorting at inference time. We fine‐tuned pretrained models for 40 epochs using an exponential temperature annealing schedule to progressively sharpen the sorting behavior. Specifically, the temperature at epoch $e$ is defined as: $\mathrm{temp}(e)\;=\;\gamma\,\mathrm{temp}(e-1),$ with $\gamma=0.8$. This decay schedule reduces the temperature to approximately $10^{-6}$, making the soft sorting operation nearly equivalent to a hard permutation. After fine‐tuning, we replaced soft sorting with hard sorting for final evaluation. Table 1 demonstrates that fine-tuning with temperature annealing and switching to hard sorting yields consistent improvements in test accuracy.
On the choice of Slicer
We employ axis-aligned slicing because our keys and queries are learned parameters; any optimal slice orientation is thus implicitly encoded in the projection matrices $W_{Q}$ and $W_{K}$. This avoids introducing extra learnable parameters and ensures a fair comparison with baselines such as vanilla attention and Sinkformer.
However, for an ablation study, we ran additional experiments on the Cats vs. Dogs dataset, varying the number of slices $L$, the inverse temperature $\tau$, and whether the slicer was learnable or frozen. Tables 6 summarizes the results. Across all settings, axis-aligned slicing provides a simple, interpretable baseline without parameter overhead. Learnable slicers can boost performance when $L$ is small, by focusing on the most informative directions, but their advantage diminishes as $L$ grows. Frozen slicers likewise require sufficiently many slices to approximate the distribution structure effectively, since performance degrades when $L$ is low in high-dimensional spaces.
@@ 5 Conclusion
We introduced ESPFormer, a novel Transformer architecture that integrates a fast, doubly-stochastic attention mechanism based on Expected Sliced Transport Plans (ESP). By leveraging differentiable soft sorting and a fully parallelizable approach, ESPFormer provides a computationally efficient alternative to iterative Sinkhorn normalization while ensuring a balanced distribution of attention across tokens. Our experiments across diverse domains, including image classification, point cloud processing, sentiment analysis, and neural machine translation, demonstrate that ESPFormer consistently outperforms both classical self-attention and Sinkhorn-based alternatives in terms of accuracy and efficiency. Furthermore, we showed that our proposed attention can be seamlessly integrated into pre-trained Transformers, improving performance even with minimal fine-tuning. The flexibility of ESPFormer also makes it compatible with emerging differential attention mechanisms, e.g., (Ye et al., 2025), expanding its applicability to a broad range of architectures. These findings highlight the potential of transport-based attention mechanisms as a principled alternative to existing methods, paving the way for future research in efficient and structured attention models.
Limitations
Our method has two main limitations. First, it is incompatible with strictly causal Transformer or autoregressive architectures, since enforcing both lower-triangularity and doubly stochasticity reduces the mapping to the identity. Second, during training its memory usage scales linearly with the number of slices, as we must retain each slice’s permutation matrix; after fine-tuning and switching to hard sorting, however, these matrices can be discarded and all inference computations can be run off-GPU.
Impact Statement
ESPFormer significantly accelerates the computation of balanced attention by introducing an efficient mechanism for enforcing doubly stochastic constraints through the mathematics of sliced optimal transport. The proposed attention improves model robustness and interoperability, and is shown to be effective even in a plug-and-play setting. This enables practical use of balanced attention as a strong inductive bias, particularly effective in low-data regimes. Furthermore, by establishing a connection between sliced optimal transport and attention mechanisms, ESPFormer opens a new research direction for structured and efficient attention through transport-based formulations. Although ESPFormer has the potential to impact a broad range of applications, we do not foresee any specific societal risks requiring special attention.
@@ Acknowledgment
This work was partially supported by the NSF CAREER Award #2339898 and the Wellcome Leap ‘Surgery: Assess/Validate/Expand (SAVE)’ program.
Appendix A Full runtime wall-clock analysis, and Patch size impact
A.1 Full runtime wall-clock
@@
A.2 Impact of Patch Size
Similar to Sander et al. (2022) and to analyze the effect of patch size on model performance, we trained Transformer, DiffTransformer, Sinkformer, and ESPFormer on the MNIST dataset (LeCun et al., 1998). To isolate the impact of the attention mechanism, we used a single-layer self-attention module without feed-forward layers. Figure 5 illustrates the test accuracy as a function of patch size for each model. ESPFormer consistently outperforms competing models, particularly for smaller patch sizes, highlighting its superior ability to capture fine-grained details. As the patch size increases, the performance gap narrows, with all models converging to similar accuracies — likely due to the reduced information content within each individual patch.
@@ C.1 Table of Notations
$\tau$ The ”inverse temperature” hyperparameter in ESP
@@ C.2 ModelNet 40
The model architecture consists of two Induced Set Attention Blocks (ISAB) in the encoder, followed by a decoder incorporating a Set Attention Block (SAB) and a Pooling by Multihead Attention (PMA) module. The training procedure employs a batch size of 64 and utilizes the Adam optimizer (Kingma & Ba, 2015). The network is trained for 300 epochs, with an initial learning rate of $10^{-3}$, which is reduced by a factor of 10 after 200 epochs.
@@ C.3 Sentiment Analysis
We also utilize this implementation for TweetEval, and similar hyperparameters for this sentiment analysis experiment can be found in table 11.
t $\tau$ $\epsilon$ $\mathcal{S}$ $\lambda_{\text{initial}}$ $lr_{initial}$ $I$
$10^{-3}$ $10^{+1}$ $10^{-1}$ 15 0.33 $10^{-4}$ ${Identity_{N\times N}}$
Table 11: Hyperparameters used in training for Transformers on TweetEval dataset.
@@ C.5.1 Cats and Dogs Classification
For this experiment, we use the ViT model (Dosovitskiy et al., 2021) with different attention mechanisms. The images are resized to $224\times 224$. We use a ViT architecture with an embedding and MLP dimension of 128, 6 layers, 8 attention heads, and a patch size of 16. For all methods and percentages of data, we train the model for 300 epochs. We use an initial learning rate of $3\times 10^{-5}$. After 250 epochs, the learning rate is reduced by a factor of 10. Training is done using the Adam optimizer (Kingma & Ba, 2015) and a batch size of 64. Our experimental setup, including the normalizations and data augmentations, are consistent with the Cats and Dogs experiment in Sinkformer (Sander et al., 2022) . For each percentage of the data, three random seeds are used to initialize and generate new subsets of data. The subsets are consistent across all methods to ensure the same training set is used.